OpenStudy (calculusxy):

ALGEBRA 1. MEDAL! Grisham is considering the three situations below. I. For the first 28 days, a sunflower grows at a rate of 3.5cm per day. II. The value of a car depreciates at a rate of 15% per year after it is purchased. III. The amount of bacteria in a culture triples every two days during an experiment. Which of the statements describes a situation with an equal difference over an equal interval?

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

They are asking which sequence of numbers is "arithmetic" as opposed to which is geometric an arithmetic sequence changes by a fixed amount. for example 1,2,3 each term is 1 more than the previous term geometric means the numbers are multiplied by a constant factor for example 1,2,4,8 each is 2 times the previous. This sequences are "exponential"

OpenStudy (calculusxy):

i don't understand

OpenStudy (phi):

if it helps, you should learn that populations grow exponentially money that earns interest grows exponentially

OpenStudy (phi):

for the sunflower, assume it starts as a seed, with zero height. can you list its height for the first few days?

OpenStudy (calculusxy):

day 1 - 0 day 2 - 3.5cm day 3 - 7 cm ...

OpenStudy (phi):

now think about " equal difference over an equal interval?" an "equal interval" could be "1 day" how much does the sunflower's height change in 1 day? 3.5 cm 0, 3.5, 7, 10.5... the difference between any pair is 3.5 cm, and the "interval" is 1 day

OpenStudy (phi):

the other choices will not have an "equal difference" (because "exponential growth" which is what they have, has an "equal multiplier"... that is different)

OpenStudy (calculusxy):

i still can't understand the difference between those two.

OpenStudy (phi):

The value of a car depreciates at a rate of 15% per year after it is purchased. that is a complicated way to say that the car starts at some price, call it P (or 100 if you don't like letters) the next year it is 15% less, P(2) = P-0.15 P or (1-0.15)P (if you know algebra) or P(2)= 0.85* P and the year 3, we would have P(3)= 0.85*0.85*P we are multiplying by 0.85 to get the next year that is different from say P(2) = P- 10 where we subtract off 10 each time.

OpenStudy (calculusxy):

so the car depreciating is exponential decay right?

OpenStudy (phi):

yes. and if its exponential (decay or growth) then it's *not* arithmetic (and arithmetic means a "constant difference" )

OpenStudy (calculusxy):

and when it is not arithmetic then it is not going in equal intervals right?

OpenStudy (phi):

yes. by "equal intervals" they mean from term 1 to term 2, term 2 to term 3, etc

OpenStudy (calculusxy):

and the fact that it's tripling every two days would not go at a constant rate right? like how it's going with the rate of 3.5cm per day.

OpenStudy (phi):

yes. if it triples every 2 days, then the sequence (starting at 1) would be 1,3,9,27... and an "equal interval" is 2 days (between numbers) and the differences are 2,6,18 clearly not equal differences

OpenStudy (calculusxy):

okay so equal differences would be the same number of difference within the same period of time (like days, weeks, etc.)

OpenStudy (phi):

yes, and if for example, you checked the sunflower case, and checked the difference between every 2 days, it would be 7 cm between any two days.

OpenStudy (calculusxy):

thanks! can u help me with a couple more of algebra 1? it's in prep for my regents.

OpenStudy (phi):

if you make a new post.

OpenStudy (calculusxy):

okay